# Extracting Sentence Embeddings from Pretrained Transformer Models

### v1

@@ 3 Methods
Here we present our investigated methods for producing sentence-level embeddings from pre-trained transformer models.
@@ 3.1 Problem formulation
and a model representing each token as a $d$-length vector $v\in\mathbb{R}^{d}$ in each of its $L$ layers. We also use static token embeddings denoted as the -1st layer and input embeddings denoted as the 0th layer. They are sums of token, segment, and position vectors, with layer normalization applied on top. Therefore, we get the following 3-dimensional representation space $\mathbb{R}^{(L+2)\times N\times d}$:
@@
We want to find such an aggregation function $f$ that the $\left((L+2)\times N\times d\right)$-dimensional output from a chosen transformer model would be reduced to as meaningful as possible $d$-dimensional vector in $\mathbb{R}^{d}$.
As a baseline for bert-base-uncased model with $L=12$ layers we consider the average over the $N$ tokens in text and over the first and last layers
@@
Generally, a corpus has $S$ number of text samples. Therefore, other text sequences can also be used to derive the representation of the target sentence. In this way, our aggregation function $f$ operates on an $\left(S\times(L+2)\times N\times d\right)$-dimensional output from the model. $S$ can also be enlarged by using additional datasets. In particular, we used Wikitext-2 from [201]. This can be used for various post-processing techniques such as centering, standardization, PCA, and others, where transformations are learned on other than the target corpus.
@@ 3.2 Models
We use multiple text representation methods, focusing mainly on BERT-based ones.
@@ 3.2.1 BERT
BERT is based on the original transformer architecture of [2]. The main difference is that instead of sequence-to-sequence workings and encoder-decoder structure, it is composed only of an encoder side. This allows it to solve multiple classification and regression tasks maintaining the same pre-trained base while only changing the classifier heads on top.
In this work, we use the BERT version called bert-base-uncased, a 110M parameter model containing 12 layers (blocks) and working with lower-cased text. We employ the Hugging Face implementation [202] of BERT. It was pre-trained on BooksCorpus (800M words) [203] and English Wikipedia (2 500M words) datasets, which take 13GB of plain text combined. The model was trained for 40 epochs (or passes through the corpus). Back then, the training time, model size, and data used were considered to be very large, yet now they are only a small fraction of what the current state-of-the-art models use.
@@
Input to the BERT encoder consists of a token, positional, and token type embeddings. BERT uses WordPiece tokenization [204]. Splitting less frequent words into sub-words (e.g., transformer $\rightarrow$ transform and ##er) rather than splitting everything on word boundaries allows one to reach a manageable vocabulary size of 30 000 tokens. BERT, like other transformer models, perceives input as a set; therefore, order information must be supplied in addition. To achieve this, additional positional embeddings are used, with a unique value for each position in a token sequence and feature dimension. Token type embeddings allow the model to distinguish between the two (if there are two) separate sentences (e.g., $\langle$ Question, Answer $\rangle$) in one token sequence. Overall, token, positional, and token type embeddings are added, and then layer normalization [205] and dropout [206] are applied. This results in the input to the BERT model blocks (for simplicity, we will call them layers).
There are 3 special tokens: [CLS] starts every token sequence, [SEP] ends token sequences and acts as a separator in a pair of sentences, and [MASK] is used during pre-training to mask some percentage of the input tokens at random for the model to predict (known as a Cloze task [207]). The [CLS] token is also used for the next sentence prediction task – predicting whether the second sentence in a pair actually follows the first one in the training dataset or is a random one – the second unsupervised pre-training task of BERT.
@@
Each transformer block (or layer) has two sublayers. The first is a multi-head self-attention mechanism and the second is a position-wise fully connected feed-forward network. The output of each sublayer Sublayer(x) is added to the input x that bypasses the sublayer through residual connection [208] and the resulting signal ends in layer normalization [205] LayerNorm(x + Sublayer(x)) becoming the input to the second sublayer or the next layer.
@@
An input tensor $X\in\mathbb{R}^{b\times t\times f}$ with $b$ samples in a batch, $t$ tokens and $f$ features is first linearly projected into queries $Q\in\mathbb{R}^{b\times t_{Q}\times h\times f_{Q}}$, keys $K\in\mathbb{R}^{b\times t_{K}\times h\times f_{K}}$, and values $V\in\mathbb{R}^{b\times t_{V}\times h\times f_{V}}$ tensors with $h$ heads using learned weights. Note that in the standard case $t=t_{Q}=t_{K}=t_{V}$ and $f=hf_{Q}=hf_{K}=hf_{V}$. Then for each sample text in a batch and each head dot products between feature vectors of every combination of query and key tokens are calculated, resulting in an attention tensor $(QK)\in\mathbb{R}^{b\times h\times t_{Q}\times t_{K}}$. It is then scaled by dividing it with $\sqrt{f_{K}}$ and softmax is applied along $t_{K}$ so that all the dot products for any given query token and all the key tokens would sum to 1. Now, a dot product is calculated again between such probability-like scores and values tensor, resulting in a weighted selection of V, with the form of $\mathbb{R}^{b\times h\times t_{Q}\times f_{V}}$. Finally, the heads are concatenated back and a linear transformation layer $W^{O}\in\mathbb{R}^{hf_{V}\times f}$ is applied. The resulting tensor is the same shape as the input one: $\mbox{MultiHead}(X)\in\mathbb{R}^{b\times t\times f}$.
For more detailed explanation of BERT and transformer architecture, please read the original papers [5, 2].
@@ 3.2.2 Prompting method (T0, T4)
Classical use of the BERT model involves two steps: 1) a general pre-training on a very large corpora with unsupervised tasks; and 2) a supervised fine-tuning step on a small target task dataset. The first step is expensive, it requires a lot of data, time, and computing resources, but it is done only once. Weights that are produced in an unsupervised way contain a lot of useful representations for the fine-tuning to only perfect them.
Prompt-based methods take advantage of the first step and can completely avoid the fine-tuning stage. The idea is to frame the target task in the original pre-training format, for which the model is essentially optimized. To extract a general sentence representation, the authors of [15] proposed to use This sentence : "[X]" means [MASK] template (which we will call T0), where the target sentence is placed instead of [X] and the final layer representation of [MASK] is used. This way the model itself tries to predict the meaning of a given sentence, and we can extract prediction weights before they are turned into probabilities over tokens.
@@
Table 1: Manual template search by adding additional text and [MASK] ([M]) tokens. Target sentences are inserted into the place of [X]. Only the average of 12th layer [M] representations are used.
@@
We construct static tokens using the Wikitext-103 [201] dataset and use the same tokenization as of BERT version bert-base-uncased. For each of all 28 807 tokens occurring in the Wikitext-103, we sum all vectors produced by the BERT model (which differ due to different contexts) and divide them by their count. Note that some tokens occur very often while others are rare (for example, the most frequent the is repeated 6 470 356 times). The process is repeated for every BERT layer, as well as combinations of layers, so that we can observe the performance dependence on this factor as well.
@@
We also wanted to see how sentence embeddings derived from static averaged BERT tokens can contribute to the original BERT representations. Therefore, we averaged sentence representations from the two methods mentioned above.
@@ 3.2.4 BERT2Static (B2S, B2S-100)
There are more advanced distillation methods than simple BERT token averaging. For example, the authors of [49] trained static word vectors using BERT’s contextualized representations. They adapted the Sent2Vec [48] model for words and trained it to predict the word given the context element of it by the contextual representation of the BERT model. This way, the authors obtained vectors for the 750 000 most frequent words. They also showed that this approach results in vectors that perform better than existing static embeddings trained from scratch, while still enjoying a small memory and computational footprint.
In our experiments, we used bert_12layer_para model [^1] Downloaded from https://github.com/epfml/X2Static provided in [49]., which was trained from the contexts of a paragraph rather than only a sentence. We evaluated two versions of BERT2Static: a regular one (B2S) and one with the 100 most frequent words filtered out (B2S-100).
@@ 3.3 Aggregating tokens
Every text contains multiple tokens, each with a corresponding embedding vector. To get one vector for the whole text out of the many, usually, a simple average is calculated. We use it as a baseline here for multiple methods.
We investigate different weighting and filtering of tokens based on their frequencies. In classical bag-of-words approaches it is usually accounted for with the tf-idf weighting. We, however, use only idf weighting, because the same tokens in BERT cannot be simply counted due to different contextualization. Given a dataset with $N$ documents and document frequency $\mbox{df}_{t}$, defined to be the number of documents in the given dataset that contain a token $t$, we calculate $\mbox{idf}_{t}$ for a token $t$ as
@@
To account for long/short text differences, we scale $\mbox{idf}_{t}$ token weights for each text so that they would sum up to 1. This way, to calculate the average embedding for a given text, one needs to only sum up the weighted vectors. We calculate two inverse document frequencies: $\mbox{idf}_{t}^{W}$ for Wikitext-2 dataset from [201] and $\mbox{idf}_{t}^{T}$ for all samples from the given target task.
We also adopt the method in [63, 15] to drop the most frequent tokens. We choose 33 tokens which were depicted in the appendix of article [15]. Following that article, we also investigate dropping punctuation and subword tokens. We found that dropping all three parts: the most frequent, punctuation, and subword tokens has the best effect, and following the original work we name such token aggregation as with removed biases (“-biases”).
@@ 3.4 Post-processing embeddings
It is common in machine learning to standardize datasets as most methods are designed to work best with normally distributed data: Gaussian with zero mean and unit variance. Yet it is not that trivial, as data may not follow a smooth distribution and may contain disturbing outliers. In the context of the best practices reviewed in Section 2.2, we investigate the following processing methods of embeddings.
z-score normalization
@@
Quantile normalization to uniform distribution (quantile-u)
This technique works by making two distributions identical in statistical properties, thus reshaping given data values according to some known distribution function. In our case, we found that the uniform distribution worked very well as a reference. For a more detailed description of the technique see [209].
We also tried other methods that are more robust to outliers, however, found their performance marginally below quantile-u. This includes quantile normalization using normal distribution and RobustScaler (as is called in the scikit-learn preprocessing library [210], which we used). The latter method removes the median value instead of the mean and scales the data according to the selected quantile range.
Whitening
It is a transformation that produces uncorrelated components each with a variance of 1.
All-but-the-top (abtt-d)
It is a method introduced in [118]. We start with the given embedding matrix $A\in\mathbb{R}^{b\times f}$ of $b$ sentences, each with $f$ features. First, it is centered by its mean $\mu\in\mathbb{R}^{f}$ into the $\tilde{A}\in\mathbb{R}^{b\times f}$. Using the centered $\tilde{A}$, and given the number $d$ of the top principal components to remove, we then calculate PCA components $U\in\mathbb{R}^{d\times f}$. Now we project our data into these components to acquire $A^{PCA}\in\mathbb{R}^{b\times f}$:
@@
The only difference of our approach from the original authors of [118] is that instead of post processing words, we use All-but-the-top to post process documents.
Normalization
@@
Learning post-processing
@@ 3.5.1 Clustering tasks
agnews 8 000 4 1 21 062 16 140 23 26 100 86
biomedical 20 000 20 1 18 888 9 326 13 20 98 64
googleTS 11 109 152 143 19 508 14 763 28 33 100 85
searchsnippets 12 340 8 7 30 643 16 334 19 24 93 77
stackoverflow 20 000 20 1 22 909 7 332 8 12 87 71
@@
It is a subset of PubMed data distributed by BioASQ [^2] http://participants-area.bioasq.org/, where 20 000 paper titles from 20 groups are randomly selected by [213].
@@
It contains titles and snippets of 11 109 news articles related to 152 events [214]. We use the full version of the dataset, which includes both titles and snippets, named GoogleNews-TS in [212].
@@
It is extracted from web search snippets and contains 12 340 snippets associated with 8 groups [215].
@@
It is a subset of the challenge data published by Kaggle [^3] https://www.kaggle.com/competitions/predict-closed-questions-on-stack-overflow/data, where [213] selected 20 000 question titles associated with 20 different categories.
@@
We perform clustering by running K-Means with the scikit-learn [216] package and reported the clustering accuracy, computed using the Hungarian algorithm [217] and averaged over 10 independent runs [^4] We used the codebase from https://github.com/amazon-science/sentence-representations and downloaded clustering datasets from https://github.com/rashadulrakib/short-text-clustering-enhancement/tree/master/data.

### v2

@@ 3 Methods
Our extensive literature review allowed us to see the big picture of the sentence embedding research.
Currently, the evolution of models for sentence embeddings and related NLP tasks is settled at transformers. In this work, we use existing models to source the raw, token-level embeddings. In Section 3.2, we describe the main model that we use in detail, some baselines we used for comparisons, as well as some of our original extensions. In particular, in T1–T4 models, we extend original prompting templates by incorporating more than one [MASK] token. Next, we present a new Avg. model where we derived sentence embeddings first by averaging tokens in different contexts and then by averaging the tokens themselves. Finally, we present our BERT + Avg. model, which combines both contextual and multiple-context averaged representations, all derived from the same BERT transformer model.
In addition to these extensions, we found two main directions that can be used to improve sentence embeddings from transformer models: token aggregation and post-processing techniques. Note, however, that our main contribution here is not the methods, as we reuse most of them from the existing works, but the combinations of them on the transformer model and extensive evaluation on multiple tasks. We thoroughly describe the techniques used (and minor extensions) for token aggregation in Section 3.3 and post-processing of embeddings in Section 3.4.
We have noticed that most works confine themselves to a small subset of evaluation tasks, which limits their results’ comparability to others. Papers from top conferences always include classification tasks on top of semantic textual similarity, which is usually the only evaluation. In this work, we evaluate sentence embeddings on three different types of tasks: semantic textual similarity, downstream classification, and clustering. We present these tasks in Section 3.5.
We now proceed with the formal definition of the problem we address in this work.
@@ 3.1 Problem Formulation
and a model representing each token as a $d$-length vector $v\in\mathbb{R}^{d}$ in each of its $L$ layers. We also use static token embeddings denoted as the -1st layer and input embeddings denoted as the 0th layer. They are sums of token, segment, and position vectors, with layer normalization applied on top. Therefore, we obtain the following 3-dimensional representation space $\mathbb{R}^{(L+2)\times N\times d}$:
@@
We want to find such an aggregation function $f$ that the $\left((L+2)\times N\times d\right)$-dimensional output from a chosen transformer model would be reduced to as meaningful a $d$-dimensional vector as possible in $\mathbb{R}^{d}$.
As a baseline for the bert-base-uncased model with $L=12$ layers, we consider the average over the $N$ tokens in text and over the first and last layers.
@@
Generally, a corpus has $S$ number of text samples. Therefore, other text sequences can also be used to derive the representation of the target sentence. In this way, our aggregation function $f$ operates on an $\left(S\times(L+2)\times N\times d\right)$-dimensional output from the model. $S$ can also be enlarged by using additional datasets. In particular, we used Wikitext-2 from [201]. This can be used for various post-processing techniques such as centering, standardization, PCA, and others, where transformations are learned on corpora other than the target corpus.
@@ 3.2 Models
We use multiple text representation methods, focusing mainly on BERT-based ones. Prompting method T4, Averaged BERT (Avg.), BERT + Avg., B2S-100, and Random embeddings (RE) are our proposed models or modifications, while BERT, T0, and B2S are plain adaptations of existing ones. We tested token aggregation and sentence representation post-processing techniques on all eight models. We will describe them in more detail below.
@@ 3.2.1 BERT
BERT is based on the original transformer architecture of [2]. The main difference is that instead of sequence-to-sequence workings and encoder–decoder structure, it is composed only of an encoder side. This allows it to solve multiple classification and regression tasks, maintaining the same pre-trained base while only changing the classifier heads on top.
In this work, we use the BERT version called bert-base-uncased, a 110 M parameter model containing 12 layers (blocks) and working with lowercase text. We employ the Hugging Face implementation [202] of BERT. It was pre-trained on BooksCorpus (800 M words) [203] and English Wikipedia (2500 M words) datasets, which take 13 GB of plain text combined. The model was trained for 40 epochs (or passes through the corpus). Back then, the training time, model size, and data used were considered to be very large, yet now they are only a small fraction of what the current state-of-the-art models use.
@@
Input to the BERT encoder consists of token, positional, and token type embeddings. BERT uses WordPiece tokenization [204]. Splitting less frequent words into sub-words (e.g., “transformer” $\rightarrow$ “transform” and “##er”) rather than splitting everything on word boundaries allows one to reach a manageable vocabulary size of 30,000 tokens. BERT, like other transformer models, perceives input as a set; therefore, order information must be supplied in addition. To achieve this, additional positional embeddings are used, with a unique value for each position in a token sequence and feature dimension. Token type embeddings allow the model to distinguish between the two (if there are two) separate sentences (e.g., $\langle$ Question, Answer $\rangle$) in one token sequence. Overall, token, positional, and token type embeddings are added, and then layer normalization [205] and dropout [206] are applied. This results in the input to the BERT model blocks (for simplicity, we will call them layers).
There are 3 special tokens. [CLS] starts every token sequence, [SEP] ends token sequences and acts as a separator in a pair of sentences, and [MASK] is used during pre-training to mask some percentage of the input tokens at random for the model to predict (known as a “Cloze” task [207]). The [CLS] token is also used for the next sentence prediction task—predicting whether the second sentence in a pair actually follows the first one in the training dataset or is a random one—the second unsupervised pre-training task of BERT.
@@
Each transformer block (or layer) has two sublayers. The first is a multi-head self-attention mechanism and the second is a position-wise fully connected feed-forward network. The output of each sublayer Sublayer(x) is added to the input x that bypasses the sublayer through residual connection [208] and the resulting signal ends in layer normalization [205], LayerNorm(x + Sublayer(x)) becoming the input to the second sublayer or the next layer.
@@
An input tensor $X\in\mathbb{R}^{b\times t\times f}$ with $b$ samples in a batch, $t$ tokens, and $f$ features is first linearly projected into queries $Q\in\mathbb{R}^{b\times t_{Q}\times h\times f_{Q}}$, keys $K\in\mathbb{R}^{b\times t_{K}\times h\times f_{K}}$, and values $V\in\mathbb{R}^{b\times t_{V}\times h\times f_{V}}$ tensors with $h$ heads using learned weights. Note that in the standard case, $t=t_{Q}=t_{K}=t_{V}$ and $f=hf_{Q}=hf_{K}=hf_{V}$. Then, for each sample text in a batch and each head dot, products between feature vectors of every combination of query and key tokens are calculated, resulting in an attention tensor $(QK)\in\mathbb{R}^{b\times h\times t_{Q}\times t_{K}}$. It is then scaled by dividing it by $\sqrt{f_{K}}$, and softmax is applied along $t_{K}$ so that all the dot products for any given query token and all the key tokens would sum to 1. Now, a dot product is calculated again for the probability-like scores and values tensor, resulting in a weighted selection of V, with the form of $\mathbb{R}^{b\times h\times t_{Q}\times f_{V}}$. Finally, the heads are concatenated back and a linear transformation layer $W^{O}\in\mathbb{R}^{hf_{V}\times f}$ is applied. The resulting tensor is the same shape as the input one: $\mbox{MultiHead}(X)\in\mathbb{R}^{b\times t\times f}$.
For a more detailed explanation of the BERT and transformer architecture, please read the original papers [5, 2].
@@ 3.2.2 Prompting Method (T0, T4)
The classical use of the BERT model involves two steps: (1) a general pre-training on a very large corpora with unsupervised tasks, and (2) a supervised fine-tuning step on a small target task dataset. The first step is expensive, it requires a lot of data, time, and computing resources, but it is carried out only once. Weights that are produced in an unsupervised way contain a lot of useful representations for the fine-tuning to only perfect them.
Prompt-based methods take advantage of the first step and can completely avoid the fine-tuning stage. The idea is to frame the target task in the original pre-training format, for which the model is essentially optimized. To extract a general sentence representation, the authors of [15] proposed using “This sentence: "[X]" means [MASK]” template (which we will call T0), where the target sentence is placed instead of [X] and the final layer representation of [MASK] is used. This way, the model itself tries to predict the meaning of a given sentence, and we can extract prediction weights before they are turned into probabilities over tokens.
@@
Table 1: Manual template search by adding additional text and [MASK] ([M]) tokens. Target sentences are inserted into the place of [X]. Only the average of 12th layer [M] representations are used. Bolded results are the best.
@@
We construct static tokens using the Wikitext-103 [201] dataset and use the same tokenization as of BERT version bert-base-uncased. For each of all 28,807 tokens occurring in the Wikitext-103, we sum all vectors produced by the BERT model (which differ due to different contexts) and divide them by their count. Note that some tokens occur very often, while others are rare (for example, the most frequent, “the”, is repeated 6,470,356 times). The process is repeated for every BERT layer, as well as combinations of layers, so that we can observe the performance dependence on this factor as well.
@@
We also wanted to see how sentence embeddings derived from static averaged BERT tokens can contribute to the original BERT representations. Therefore, we averaged sentence embeddings from the two methods mentioned above.
@@ 3.2.4 BERT2Static (B2S, B2S-100)
There are more advanced distillation methods than simple BERT token averaging. For example, the authors of [49] trained static word vectors using BERT’s contextualized representations. They adapted the Sent2Vec [48] model for words and trained it to predict the word given the context element of it by the contextual representation of the BERT model. This way, the authors obtained vectors for the 750,000 most frequent words. They also showed that this approach results in vectors that perform better than existing static embeddings trained from scratch, while still enjoying a small memory and computational footprint.
In our experiments, we used the bert_12layer_para model (downloaded from https://github.com/epfml/X2Static, provided in [49], accessed on 1 May 2023), which was trained from the contexts of a paragraph rather than only a sentence. We evaluated two versions of BERT2Static: a regular one (B2S) and one with the 100 most frequent words filtered out (B2S-100).
@@ 3.3 Aggregating Tokens
Every text contains multiple tokens, each with a corresponding embedding vector. To obtain one vector for the whole text out of the many, usually, a simple average is calculated, as discussed in Section 2.1. We use it as a baseline here for multiple methods.
We investigate different methods of token weighting and filtering based on their frequencies. In classical bag-of-words approaches, it is usually accounted for with the tf-idf weighting. We, however, use only idf weighting, because the same tokens in BERT cannot be simply counted due to different contextualization. Given a dataset with $N$ documents and document frequency $\mbox{df}_{t}$, defined to be the number of documents in the given dataset that contain a token $t$, we calculate $\mbox{idf}_{t}$ for a token $t$ as
@@
To account for long/short text differences, we scale $\mbox{idf}_{t}$ token weights for each text so that they would sum up to 1. This way, to calculate the average embedding for a given text, one need only sum up the weighted vectors. We calculate two inverse document frequencies: $\mbox{idf}_{t}^{W}$ for the Wikitext-2 dataset from [201] and $\mbox{idf}_{t}^{T}$ for all samples from the given target task.
We also adopt the method in [63, 15] to drop the most frequent tokens. We choose 33 tokens, which were depicted in the appendix of [15]. Following that article, we also investigate dropping punctuation and subword tokens. We found that dropping all three parts—the most frequent, punctuation, and subword tokens—has the best effect, and following the original work, we name such token aggregation as with removed biases (“-biases”).
@@ 3.4 Post-Processing Embeddings
It is common in machine learning to standardize datasets, as most methods are designed to work best with normally distributed data: Gaussian with zero mean and unit variance. Yet it is not that trivial, as data may not follow a smooth distribution and may contain disturbing outliers. In the context of the best practices reviewed in Section 2.2, we investigate the following processing methods of embeddings.
3.4.1 Z-Score Normalization
@@
3.4.2 Quantile Normalization to Uniform Distribution (quantile-u)
This technique works by making two distributions identical in statistical properties, thus reshaping given data values according to some known distribution function. In our case, we found that the uniform distribution worked very well as a reference. For a more detailed description of the technique, see [209].
We also tried other methods that are more robust to outliers; however, we found their performance marginally below quantile-u. This includes quantile normalization using a normal distribution and RobustScaler (as it is called in the scikit-learn preprocessing library [210], which we used). The latter method removes the median value instead of the mean and scales the data according to the selected quantile range.
3.4.3 Whitening
It is a transformation that produces uncorrelated components, each with a variance of 1.
3.4.4 All-But-The-Top (ABTT)
It is a method introduced in [118]. We start with the given embedding matrix $A\in\mathbb{R}^{b\times f}$ of $b$ sentences, each with $f$ features. First, it is centered by its mean $\mu\in\mathbb{R}^{f}$ into $\tilde{A}\in\mathbb{R}^{b\times f}$. Using the centered $\tilde{A}$, and given the number $d$ of the top principal components to remove, we then calculate PCA components $U\in\mathbb{R}^{d\times f}$. Now, we project our data into these components to acquire $A^{PCA}\in\mathbb{R}^{b\times f}$:
@@ 3.4.4 All-But-The-Top (ABTT)
The only difference of our approach from the original authors of [118] is that instead of post processing words, we use all-but-the-top to post-process documents.
3.4.5 Normalization
@@
3.4.6 Learning Post-Processing
@@ 3.5.1 Clustering Tasks
agnews 8000 4 1 21,062 16,140 23 26 100 86
biomedical 20,000 20 1 18,888 9326 13 20 98 64
googleTS 11,109 152 143 19,508 14,763 28 33 100 85
searchsnippets 12,340 8 7 30,643 16,334 19 24 93 77
stackoverflow 20,000 20 1 22,909 7332 8 12 87 71
@@
It is a subset of PubMed data distributed by BioASQ ( http://participants-area.bioasq.org/, accessed on 1 May 2023), where 20,000 paper titles from 20 groups are randomly selected by [213].
@@
It contains titles and snippets of 11,109 news articles related to 152 events [214]. We use the full version of the dataset, which includes both titles and snippets, named GoogleNews-TS in [212].
@@
It is extracted from web search snippets and contains 12,340 snippets associated with 8 groups [215].
@@
It is a subset of the challenge data published by Kaggle (https://www.kaggle.com/competitions/predict-closed-questions-on-stack-overflow/data, accessed on 1 May 2023), where [213] selected 20,000 question titles associated with 20 different categories.
@@
We perform clustering by running $k$-means with the scikit-learn [216] package and reported the clustering accuracy, computed using the Hungarian algorithm [217] and averaged over 10 independent runs (we used the codebase from https://github.com/amazon-science/sentence-representations and downloaded clustering datasets from https://github.com/rashadulrakib/short-text-clustering-enhancement/tree/master/data, both were accessed on 1 May 2023).